# The Gravitational Wave Memory from Binary Neutron Star Mergers

### v1

@@ Abstract
Many studies have been performed to numerically compute the memory effect Pollney and Reisswig (2011); Mitman et al. (2020, 2021); Moxon et al. (2020, 2023); Yoo et al. (2023); Chen et al. (2024). Most of these focused on binary black holes (BBHs) that are expected to have the highest GW luminosities Lasky et al. (2016); Yang and Martynov (2018a); Boersma et al. (2020); Grant and Nichols (2023). On the other hand, binary neutron stars (BNSs) will have smaller GW luminosities, but they will exhibit additional memory contributions due to the EM and neutrino luminosities, as well as the baryonic ejecta, all of which are absent in BBH mergers. This includes contributions arising from relativistic jets and (if produced) short $\gamma$-ray bursts (sGRBs) Sago et al. (2004); Birnholtz and Piran (2013); Sakai et al. (2025). In Tiwari et al. (2021); Lopez et al. (2024) it is argued that the memory can be used to distinguish BNS from BBH mergers, while in Yang and Martynov (2018b) it is argued that it can be used to constrain the yet unknown neutron star (NS) equation of state (EOS). While Karlson (2018); Yang and Martynov (2018b); Tiwari et al. (2021); Lopez et al. (2024) investigated the memory from BNS mergers, they only computed the GW null memory. The purpose of this work is to complete the picture by quantifying—for the first time—the additional contributions to the displacement memory from the EM field, neutrinos, and ejecta using GRMHD simulations.
@@
Simulations and Methods.—Two sets of simulations are used in this work. The first uses our well-tested IllinoisGRMHD code, includes magnetic fields, but has no neutrinos, and is described in detail in Bamber et al. (2025); Tsokaros et al. (2025). The second set uses the IllinoisGRMHD thorn Etienne et al. (2015) in the Einstein Toolkit Loffler et al. (2012), and includes both magnetic fields and neutrino radiation computed via a leakage scheme (see Supplement for details).
To leading order, the ordinary GW memory in the transverse-traceless (TT) gauge for a system of $N$ bodies with masses $M_{A}$ and velocities $\bm{v}_{A}$, is given by Braginsky and Thorne (1987)
$\Delta h^{\rm ord}_{ij}=\frac{4}{r}\;\Delta\sum^{N}_{A=1}E_{A}\left[\frac{v^{i}_{A}v^{j}_{A}}{(1-\bm{v}_{A}\cdot\bm{N})}\right]^{\mathrm{TT}},$ (1)
where $\Delta$ means the difference between the initial and final states of the summation, $E_{A}=M_{A}\gamma_{A}=M_{A}/\sqrt{1-v^{2}_{A}}$ is the energy of each particle, $\bm{N}$ is a unit vector that points from the source to the observer, $r$ is the coordinate distance to the source and “TT” denotes the transverse-traceless part. Units with $G=c=1$ are used in this work unless stated explicitly. We use this formula to compute the ordinary memory due to the ejected baryonic mass which we term the “ejecta memory” (see Supplement). We note here that the ordinary memory is a subset of the linear memory and it was the memory calculated originally by Zel’dovich and Polnarev Zel’dovich and Polnarev (1974).
@@
With regards to point (3), the top and middle panels in Fig. 2 show two binaries with the same mass but different EOS whose memory content behaves differently with respect to the corresponding nonmagnetized runs. In the top panel, all EM fields but the strongest one result in very small differences in the memory relative to the nonmagnetized run. In particular, there is a small monotonic decrease in the GW null memory strain for the pulsar-like topology as the magnetic field is increased. However, for the same mass with a different EOS (ALF2, middle panel) the GW null memory of the case with smallest non-zero magnetic field $|B|_{\mathrm{max,ins}}=5.5\times 10^{15}$G (blue) is larger than that of the case with no magnetic field (black). Increasing further the magnetic field leads to a monotonic decrease of the GW null memory. This suggests that even moderate magnetic fields affect the GW null memory of BNS with identical masses differently, depending on the EOS. We also note here that the expectation that magnetized binaries will produce larger total memory signals than nonmagnetized ones due to the additional EM null memory Bieri et al. (2012) is not necessarily realized, as the EM null memory is negligible except for the highest magnetic field case considered (red).
The middle and bottom panels of Fig. 2 show results from binaries with the same ALF2 EOS but different masses. In the middle panel, the case with $|B|_{\mathrm{max,ins}}=5.5\times 10^{15}$G (blue) has the largest GW null memory, while in the bottom panel it is the case with $|B|_{\mathrm{max,ins}}=5.5\times 10^{16}$G and an interior-only toroidal topology (purple). No monotonic behavior with respect to the magnitude of the magnetic field is observed here. Figure 2 is consistent with the GW luminosity plots (see Fig. 21 in Bamber et al. (2025) as well as the $\dot{h}_{\ell m}$ plots in the Supplement).
@@
In Fig. 3 we show the dominant $l=2,m=0$ mode of the memory (GW, EM, ejecta, and neutrino) from the second set of simulations which include interior only EM fields and neutrino radiation. As in the majority of the first set of simulations, the memory is dominated by the GW null contribution, which depends strongly on the EOS (see bottom two panels) and which can either be larger or smaller than the corresponding nonmagnetized case, depending on the inserted field strength and topology. Although the seed magnetic fields are large (see Supplement for details), the EM null memory and ejecta memory are negligible, likely because the magnetic field is confined to the interior of the NS.
@@
Figure 4: Characteristic strain curves for the oscillatory (dashed) and total memory (solid) parts of the signal for selected binaries with $M=2.57M_{\odot}$ and $q=1$ at a luminosity distance of 2020 Mpc (c.f., top two panels in Fig. 2). We show zero magnetic field cases for the ALF2 and SLy EOS, and two nonzero (initially pulsar-like) magnetic field cases for ALF2.
We also show the memory for an equal-mass BBH merger with the same mass and distance. Sensitivity curves correspond to Advanced LIGO Buikema et al. (2020), the Einstein Telescope (D configuration) Hild et al. (2011), and Cosmic Explorer Reitze et al. (2019); Evans et al. (2018).
While the memory signal is roughly two orders of magnitude smaller than the oscillatory component at the same frequency, it may still be detectable with next-generation GW observatories like the Einstein Telescope and Cosmic Explorer, with a SNR of up to $\sim 5$ for sources at 50 Mpc. The mismatch between the memory signal using Cosmic Explorer for the two cases with the same mass $M=2.57M_{\odot}$, no magnetic fields, and two different EOSs (SLy and ALF2) is $0.8\%$, similar to the mismatches observed between different EOSs in Table I of Lopez et al. (2024).Applying the waveform distinguishability criterion (e.g., Eq. (44) of Bamber et al. (2025)) suggests that $\sim 30\,{\rm Mpc}$ is the maximum distance at which those two cases can be distinguished.
@@
Conclusions.—We performed the first memory computations for BNS mergers that include not only the GW null memory, but the EM, neutrino, and ejected matter memory contributions as well. EM fields can either increase or decrease the memory, depending on the field strength, its topology, EOS, and binary mass. But the EM null memory is subdominant to the GW null memory for all but the largest EM fields. The memory from the ejected matter and neutrinos is also subdominant, but can be nonnegligble for certain EOS, masses, and field strengths. Future analyses of BNS memory should also consider degeneracies between different physical phenomena, including the magnetic field and the EOS Tsokaros et al. (2025).
Movies and additional visualizations highlighting our results can be found at web (2025).

### v2

@@ Abstract
Many studies have been performed to numerically compute the memory effect Pollney and Reisswig (2011); Mitman et al. (2020); Mitman and others (2021); Moxon et al. (2020, 2023); Yoo et al. (2023); Chen and others (2024). Most of these focused on binary black holes (BBHs) that are expected to have the highest GW luminosities Lasky et al. (2016); Yang and Martynov (2018b); Boersma et al. (2020); Grant and Nichols (2023). On the other hand, binary neutron stars (BNSs) will have smaller GW luminosities, but they will exhibit additional memory contributions due to the EM and neutrino luminosities, as well as the baryonic ejecta, all of which are absent in BBH mergers. This includes contributions arising from relativistic jets and (if produced) short $\gamma$-ray bursts (sGRB) Sago et al. (2004); Birnholtz and Piran (2013); Sakai et al. (2025). In Tiwari et al. (2021); Lopez et al. (2024) it is argued that the memory can be used to distinguish BNS from BBH mergers, while in Yang and Martynov (2018a) it is argued that it can be used to constrain the yet unknown neutron star (NS) equation of state (EOS). While Karlson (2018); Yang and Martynov (2018a); Tiwari et al. (2021); Lopez et al. (2024) investigated the memory from BNS mergers, they only computed the GW null memory. The purpose of this work is to complete the picture by quantifying—for the first time—the additional contributions to the displacement memory from the EM field, neutrinos, and ejecta using GRMHD simulations.
@@
Simulations and Methods.—Two sets of simulations are used in this work. The first uses our well-tested IllinoisGRMHD code, includes magnetic fields, but has no neutrinos, and is described in detail in Bamber et al. (2025); Tsokaros et al. (2025). The second set uses the IllinoisGRMHD thorn Etienne et al. (2015) in the Einstein Toolkit Loffler and others (2012), and includes both magnetic fields and neutrino radiation computed via a leakage scheme.
To leading order, the ordinary GW memory in the transverse-traceless (TT) gauge for a system of $N$ bodies with masses $M_{A}$ and velocities $\boldsymbol{v}_{A}$, is given by Braginsky and Thorne (1987)
$\Delta h^{\rm ord}_{ij}=\frac{4}{r}\;\Delta\sum^{N}_{A=1}E_{A}\left[\frac{v^{i}_{A}v^{j}_{A}}{(1-\boldsymbol{v}_{A}\cdot\boldsymbol{N})}\right]^{\mathrm{TT}},$ (1)
where $\Delta$ means the difference between the initial and final states of the summation, $E_{A}=M_{A}\gamma_{A}=M_{A}/\sqrt{1-v^{2}_{A}}$ is the energy of each particle, $\boldsymbol{N}$ is a unit vector that points from the source to the observer, $r$ is the coordinate distance to the source and “TT” denotes the transverse-traceless part. Units with $G=c=1$ are used in this work unless stated explicitly. We use this formula to compute the ordinary memory due to the ejected baryonic mass which we term the “ejecta memory” (see Supplemental Material). We note here that the ordinary memory is a subset of the linear memory and it was the memory calculated originally by Zel’dovich and Polnarev Zel’dovich and Polnarev (1974).
@@
With regards to point (3), the top and middle panels in Fig. 2 show two binaries with the same mass but different EOS whose memory content behaves differently with respect to the corresponding nonmagnetized runs. In the top panel, all EM fields but the strongest one result in very small differences in the memory relative to the nonmagnetized run. In particular, there is a small monotonic decrease in the GW null memory strain for the pulsar-like topology as the magnetic field is increased. However, for the same mass with a different EOS (ALF2, middle panel) the GW null memory of the case with smallest non-zero magnetic field $|B|_{\mathrm{max,ins}}=5.5\times 10^{15}$G (blue) is larger than that of the case with no magnetic field (black). Increasing further the magnetic field leads to a monotonic decrease of the GW null memory. This suggests that even moderate magnetic fields affect the GW null memory of BNS with identical masses differently, depending on the EOS. We also note here that the expectation that magnetized binaries will produce larger total memory signals than nonmagnetized ones due to the additional EM null memory Bieri et al. (2012), is not necessarily realized, since on one hand the EM null memory is negligible (except for the highest magnetic field case considered here (red)), and on the other, because the postmerger magnetized evolution of the remnant NS can result in a smaller GW null memory.
The middle and bottom panels of Fig. 2 show results from binaries with the same ALF2 EOS but different masses. In the middle panel, the case with $|B|_{\mathrm{max,ins}}=5.5\times 10^{15}$G (blue) has the largest GW null memory, while in the bottom panel it is the case with $|B|_{\mathrm{max,ins}}=5.5\times 10^{16}$G and an interior-only toroidal topology (purple). No monotonic behavior with respect to the magnitude of the magnetic field is observed here. Figure 2 is consistent with the GW luminosity plots (see Fig. 21 in Bamber et al. (2025) as well as Fig. 5).
@@
In Fig. 3 we show the dominant $l=2,m=0$ mode of the memory (GW, EM, ejecta, and neutrino) from the second set of simulations which include interior only EM fields and neutrino radiation. As in the majority of the first set of simulations, the memory is dominated by the GW null contribution, which depends strongly on the EOS (see bottom two panels) and which can either be larger or smaller than the corresponding nonmagnetized case, depending on the inserted field strength and topology. Although the seed magnetic fields are large, the EM null memory and ejecta memory are negligible, likely because the magnetic field is confined to the interior of the NS.
@@
Figure 4: Characteristic strain curves for the oscillatory (dashed) and total memory (solid) parts of the signal for selected binaries with $M=2.57M_{\odot}$ and $q=1$ at a luminosity distance of 5050 Mpc (c.f., top two panels in Fig. 2). We show zero magnetic field cases for the ALF2 and SLy EOS, and two nonzero (initially pulsar-like) magnetic field cases for ALF2.
We also show the memory for an equal-mass BBH merger with the same mass and distance. Sensitivity curves correspond to Advanced LIGO Buikema and others (2020), the Einstein Telescope (D configuration) Hild and others (2011), and Cosmic Explorer Reitze and others (2019); Evans et al. (2018).
While the memory signal is roughly two orders of magnitude smaller than the oscillatory component at the same frequency, it may still be detectable with next-generation GW observatories like the Einstein Telescope and Cosmic Explorer, with a SNR of up to $\sim 5$ for sources at 50 Mpc. Following Lopez et al. (2024), we can test whether the memory signal alone (i.e. excluding the oscillatory part) can be used to constrain the NS EOS. The mismatch between the memory signals using Cosmic Explorer for the two cases with the same mass $M=2.57M_{\odot}$, no magnetic fields, and two different EOSs (SLy and ALF2) is $0.8\%$, similar to the mismatches observed between different EOSs in Table I of Lopez et al. (2024). Applying the waveform distinguishability criterion (e.g., Eq. (44) of Bamber et al. (2025)) suggests that $\sim 30\,{\rm Mpc}$ is the maximum distance at which those two cases can be distinguished.
@@
Conclusions.—We performed the first memory computations for BNS mergers that include not only the GW null memory, but the EM, neutrino, and ejected matter memory contributions as well. EM fields can either increase or decrease the memory, depending on the field strength, its topology, EOS, and binary mass. But the EM null memory is subdominant to the GW null memory for all but the largest EM fields. The memory from the ejected matter and neutrinos is also subdominant, but can be nonnegligble for certain EOS, masses, and field strengths. Future analyses of BNS memory should also consider degeneracies between different physical phenomena, including the magnetic field and the EOS Tsokaros et al. (2025). Movies and additional visualizations highlighting our results can be found at 1.